OpenStudy (anonymous):

Write the equation of the line which passes through (5, –2) and is parallel to x = 4.

OpenStudy (ivettef365):

use formula y=mx+b

OpenStudy (anonymous):

wud -2 be the b?

OpenStudy (anonymous):

and 5 be mx

OpenStudy (ivettef365):

you have point (5, -2) and slope = 4 then to find b substitute the points y = mx + b -2=(4)(5) + b -2 = 20 + b can you tell me what is b

OpenStudy (ivettef365):

Now just make sure when you said x = 4 are you saying m = 4, or slope = 4

OpenStudy (anonymous):

wub b = -22

OpenStudy (ivettef365):

correct so now you have your m and your b that is what you need to write the equation of the line y = 4x -22

OpenStudy (anonymous):

so y=4x-22 thts my finished equation?

OpenStudy (anonymous):

Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.

OpenStudy (anonymous):

this is a different one wud i still be using y =mx+b

OpenStudy (ivettef365):

to be perpendicular the slope needs to be opposite reciprocal, what is the opposite reciprocal of -2 ??

OpenStudy (anonymous):

positive 2

OpenStudy (ivettef365):

opposite is positive 2, but opposite reciprocal = positive 1/2

OpenStudy (ivettef365):

so the slope = 1/2, now find b using the formula y = mx + b, where m = 1/2, and (x,y) = points (2,1)

OpenStudy (anonymous):

so if m was 2 it wud be 1? Since half

OpenStudy (ivettef365):

if m would have been 2, the opposite reciprocal would have been -1/2

OpenStudy (anonymous):

negative 2?

OpenStudy (ivettef365):

you have y = mx + b, so you have 1 = (1/2)(2) + b 1 = 1 + b, so b = 1-1, b = 0

OpenStudy (ivettef365):

so the equation of the line is written like this y = 1/2x

OpenStudy (anonymous):

but wudnt there be more to the equation then just that?

OpenStudy (ivettef365):

when a question ask you "What is the equation of the line", is asking to put the equation on the form y = mx + b all you need on this equation is the value of m and b, just because (x,y) represents any points on that line, and a line have many, many points.

OpenStudy (anonymous):

oh alright math is my worst subject -_- its confusing me takes me a lil to catch on

OpenStudy (ivettef365):

that is ok, math is all about practice

OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. Now this one is the slope intercept

OpenStudy (ivettef365):

the y=mx+b, is the equation of the line in slope intercept form

OpenStudy (ivettef365):

so you have y = -x+6 what is the slope or m on this equation ??

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

when u have (-4,2) wud the first one represent m or the second one

OpenStudy (ivettef365):

the points are (-4,2), that have nothing to do with slope on the equation y = mx + b y is a point x is a point m is the SLOPE b is the y - int

OpenStudy (anonymous):

ok so how do u find slope

OpenStudy (ivettef365):

ok, is giving you the equation y = -x + 6, what number is on the m position

OpenStudy (anonymous):

-x

OpenStudy (ivettef365):

right but x is a point (-) is the slope which is the same as -1

OpenStudy (ivettef365):

now that you now the slope = -1, to find the equation of the line which passes through (–4, 2) and is parallel to the equation when they ask for "PARALLEL" it means the slope would be the same

OpenStudy (ivettef365):

so all you need to find now is the value of b, so you have M = -1 X = -4 Y = 2 equation y = mx + b, what is the value of b

OpenStudy (anonymous):

ok i got 2=-1(4) tht became -4 but still tht didnt give b

OpenStudy (anonymous):

so then i wud have 2=-4

OpenStudy (anonymous):

wud i divide 2 on each side to find b

OpenStudy (ivettef365):

you have 2 = (-1)(-4) + b

OpenStudy (ivettef365):

solve for b

OpenStudy (anonymous):

6=b

OpenStudy (ivettef365):

2 = 4 + b -4 -4 ---------- -2 = b

OpenStudy (ivettef365):

b = -2

OpenStudy (ivettef365):

so now you have the m and the b m = -1 b = -2

OpenStudy (anonymous):

how in the world did i get 6 -_-

OpenStudy (ivettef365):

so what is the equation

OpenStudy (anonymous):

y=-1+(-2)

OpenStudy (anonymous):

or wud it just be y=-1-2

OpenStudy (anonymous):

-1+2*

OpenStudy (ivettef365):

remember on the equation you always need to have an (X) and a (Y) so you write it as y = -x -2 or y = -1x - 2

OpenStudy (anonymous):

ohhh okkk so cant forget the "x"

OpenStudy (ivettef365):

think you are a little confuse, you don't need a value for y, you only need the "equation of the line" which is a formula written as y = mx + b, where m = slope and b = y-int

OpenStudy (anonymous):

ohhh okkkk alright with math its remembering all these formulas haha

OpenStudy (ivettef365):

I know it is very confusing, but practice practice practice

OpenStudy (ivettef365):

got to go

OpenStudy (anonymous):

Thank u!!! :)